# Traffic Co-Simulation Framework Empowered by Infrastructure Camera Sensing and Reinforcement Learning

### v1

@@ Abstract
Traffic simulations are commonly used to optimize traffic flow, with reinforcement learning (RL) showing promising potential for automated traffic signal control. Multi-agent reinforcement learning (MARL) is particularly effective for learning control strategies for traffic lights in a network using iterative simulations. However, existing methods often assume perfect vehicle detection, which overlooks real-world limitations related to infrastructure availability and sensor reliability. This study proposes a co-simulation framework integrating CARLA and SUMO, which combines high-fidelity 3D modeling with large-scale traffic flow simulation. Cameras mounted on traffic light poles within the CARLA environment use a YOLO-based computer vision system to detect and count vehicles, providing real-time traffic data as input for adaptive signal control in SUMO. MARL agents, trained with four different reward structures, leverage this visual feedback to optimize signal timings and improve network-wide traffic flow. Experiments in the test-bed demonstrate the effectiveness of the proposed MARL approach in enhancing traffic conditions using real-time camera-based detection. The framework also evaluates the robustness of MARL under faulty or sparse sensing and compares the performance of YOLOv5 and YOLOv8 for vehicle detection. Results show that while better accuracy improves performance, MARL agents can still achieve significant improvements with imperfect detection, demonstrating adaptability for real-world scenarios.
Index Terms:
@@
Computer Vision, Co-simulation Framework, Infrastructure Camera, Reinforcement Learning, Traffic Signal Control
I Introduction
@@
Traffic congestion is a significant issue in urban areas leading to increased travel times, fuel consumption, and environmental pollution, causing considerable economic impact [1, 2, 3]. Effective traffic signal control (TSC) can alleviate these problems by optimizing the flow of vehicles through intersections. Traditional TSC methods often rely on fixed-time schedules or simple rule-based systems, which may not adapt well to dynamic and complex traffic conditions [4]. As a result, there is growing interest in applying advanced techniques, such as reinforcement learning (RL), to develop adaptive and efficient traffic signal control systems.
@@
Evaluating traffic control strategies in simulation before real-world deployment is essential for intelligent transportation systems (ITS) [5]. Among the available simulation tools, Simulation of Urban MObility (SUMO) [6] stands out for its open-source availability, ease of use, and effectiveness [7]. While this is a sufficient and widely used environment for traffic signal control experimentation, the lack of realistic synthetic sensor data limits the applicability to real-world situations. Despite the widespread use of video cameras due to their low cost and ability to capture rich data [8], previous studies have yet to fully integrate them into SUMO-based traffic control systems. To bridge this gap, co-simulation frameworks that combine multiple simulators are gaining popularity as a means to create more realistic and comprehensive evaluation environments that include virtual cameras [9]. However, further work is needed to transform these cameras into smart sensors that can interact with microsimulation control algorithms in a synchronized way. This integration would create a virtual test-bed with significant potential to simulate surveillance camera-based ITS control strategies.
@@
In this study, we propose and evaluate such a promising test-bed by exploring the use of multi-agent reinforcement learning (MARL) for traffic signal optimization and using camera-based detection for testing. In order to examine the effectiveness of the co-simulation framework, we develop and evaluate MARL agents in this realistic test-bed that integrates the CARLA simulator [10] and the SUMO traffic simulator. CARLA is an open-source urban driving simulator that provides high-fidelity 3D environments for vehicle dynamics and sensor simulations, while SUMO is an open-source, highly portable, microscopic and continuous multi-modal traffic simulation package designed to handle large road networks [11].
@@
MARL involves training multiple RL agents, each controlling a traffic signal, to collectively improve overall travel time in the road network. This approach allows each traffic signal to adapt its timings based on real-time traffic conditions, leading to more efficient traffic management [12]. There are numerous ways to train these agents in terms of architecture and reward functions. However, the challenges of real-world applicability stemming from faulty or sparse sensing, communication delays, and computational bottlenecks can significantly impact the effectiveness of any control system [13], yet these issues are rarely addressed in simulation studies.
@@
In our framework, CARLA is used to provide a realistic urban environment and detailed vehicle dynamics, while SUMO handles traffic flow and signal control. Cameras are deployed on traffic light poles within the CARLA environment to monitor traffic at intersections. These cameras feed a convolutional neural network based computer vision system to detect and count vehicles, providing real-time data on traffic conditions which is used by the MARL agents to control the next phase of the traffic light. Figure 1 showcases the framework for the system. The agents have been trained using a Q-learning algorithm to optimize signal timings based on the observed traffic conditions. The goal is to minimize network-wide vehicle waiting times and improve overall traffic flow.
@@
This paper presents the methodology and results of our co-simulation study, demonstrating the potential of MARL for adaptive traffic signal control using computer vision based detection that is often imperfect. We discuss the relevant literature followed by the setup of the CARLA-SUMO co-simulation environment, the implementation of the computer vision system for vehicle detection, and the training and evaluation of the MARL agents. Our results show that MARL-based traffic signal control can significantly improve traffic flow compared to traditional fixed-time signal control methods while using inputs from infrastructure camera-based vehicle detection. We also discuss the errors related to detection and its effects on the system performance. In summary, the primary contributions of the paper are as follows:
@@
  • The framework deploys cameras on traffic light poles within the CARLA environment. These cameras use a YOLO-based computer vision system to detect and count vehicles, providing real-time data on traffic conditions in the virtual test-bed.
  • The study employs MARL with four different rewards to train multiple agents for adaptive traffic signal control, optimizing the timings based on visual feedback of real-time traffic conditions to improve network-wide traffic flow. Experiments on the co-simulation test-bed show that the proposed MARL significantly improves traffic signal control using camera-based detection.
  • The MARL agents are in scenarios with potentially faulty or sparse sensing, highlighting the importance of robustness in real-world applications. The paper examines performance differences between YOLOv5 and the newer YOLOv8, and how they affect the agents trained using different rewards, finding that while increased accuracy improves performance, even imperfect detection can yield positive results with the right MARL training.
@@ II Literature Review
There are numerous traffic simulation software available for different purposes including traffic flow modeling, simulating connected autonomous vehicles, vehicle dynamics and emissions modeling, and traffic signal control [14]. Among the most widely used traffic simulators are AIMSUN [15], VISSIM [16], and MATSim [17], each offering unique features and capabilities to suit different research needs [18]. However, the focus of this study is on SUMO [6], a highly regarded open-source traffic simulator. The open-source nature of SUMO and the Traffic Control Interface (TraCI) enable finely customizable real-time programmable interaction with the simulation. It has a rich history of use in intelligent traffic signal control [19][20].
@@
While automated traffic signal control has been around for decades, most notably the Sydney coordinated adaptive traffic system (SCATS) [21] and split, cycle, and offset optimisation technique (SCOOT) [22]. These methods utilized real-time traffic flow data at each intersection to manage the timing of signal phases. The SCATS system uses vehicle counts at one intersection whereas SCOOT relies on upstream detectors for more detailed flow information. Both systems are capable of coordination with neighboring intersections by communicating their state information. However, the systems operate in a local context, their behavior is limited by the preprogrammed algorithm, and they are sensitive to the position and reliability of the vehicle sensors [23].
@@
Machine learning strategies, in particular reinforcement learning (RL) has emerged as an effective strategy for automated traffic control [24]. Coordinated deep Q-network RL was used with transfer planning for fast and scalable learning, but the training was found to occasionally lack in stability [25]. Reinforced Signal Control (RESCO) [26] provided a benchmark test-bed for RL based TSC inspired by real world cities and calibrated demand. Their extensive experimentation showed that decentralized control algorithms are more robust as compared to coordinated control methods, which are hampered by stricter sensing requirements and do not generalize well to all traffic conditions. A scalable multi-agent RL (MARL) algorithm for TSC was developed in [27] which features independent advantage actor critic RL agents, each of which is kept updated about its neighbor’s policy. It was tested on a synthetic traffic grid and Monaco city traffic network, with the salient limitations being a lack of robustness to noisy or delayed sensor measurements, high memory consumption of the agent models, and communication latency between neighboring intersections.
@@
Regarding the interpretability of intelligent traffic control agents, it is demonstrated in [28] that a regulatable control policy is one in which changes to the signal assignment can be understood as a direct consequence of changes in the state of the traffic. This is important for safety and regulatory considerations and is comparable to the actuated signal technology currently deployed based on government standards [29].
TABLE I: Existing literature related to traffic co-simulation
Paper Task Key Method Infrastructure Sensing Scale Feedback Loop
Osiński et. al. [30] RL for autonomous driving Image segmentation + CNN No Single vehicle Yes
Cantas et. al. [31] Collaborative perception for CAVs LiDAR 3D detection + Sensor fusion No Multiple vehicles Yes
Aoki et. al. [32] Cooperative perception and V2X communication Object classification + Deep-Q RL No Multiple vehicles No
Azfar et. al. [33] Incorporating vehicle detection in Digital twin Edge-based vehicle detection Yes Single roundabout No
Shi et. al. [34] Data-driven test-bed for traffic scenes Vehicle detection and trajectory extraction Yes Highway segment or intersection No
Zhou et. al. [35] CV based traffic modeling and adaptive signal schemes Transfer learning for vehicle detection and tracking Yes Road network No
Kumaran et. al. [36] CV based traffic signal optimization Optical flow and clustering + Gaussian regression Yes Single intersection No
He et. al. [9] CV based isolated TSC End-to-end deep RL from images Yes Single intersection Yes
Zhao et. al. [37] Modeling and evaluating CAV and human behavior Mixed traffic modeling and game theory No Highway on-ramp Yes
Insuasty et. al. [38] Control strategies for urban traffic with co-simulation Particle swarm optimization + Replicator dynamics No Two intersections Yes
Zhu et. al. [39] Mixed traffic stability analysis in co-simulation Mathematical modeling with V2V and car following No Multiple vehicles No
Ours Test-bed for multi-agent traffic signal optimization using infrastructure camera input Vehicle detection + MARL Yes Multiple intersections Yes
@@
A traffic simulator such as SUMO can perform its role effectively but is limited in its ability to provide synthetic sensor data beyond loop detectors. Additional simulation tools are required for problems that require real-world considerations such as demand prediction, advanced sensing, visual accuracy, vehicle mechanics, etc. These simulations must be synchronized for consistency and validity, and this process is known as co-simulation. MATSim was integrated with external simulators for fleet demand simulation in [40] to showcase ride-pooling services with agent based simulation. CARLA is an open source simulator for autonomous driving built atop the Unreal 4 game engine [10]. It features high quality 3D graphics, a flexible API, and integration with SUMO. CARLA-SUMO co-simulation is extensively used in many applications including the study of connected and autonomous vehicles [41], real-time sensor fusion [31], computer vision based cooperative perception [32], and transportation digital twins [42]. The co-simulation has also been used in conjunction with image processing and object detection to estimate traffic flow statistics such as volume and trajectories from real world video in order to generate simulation scenarios based on real-world conditions [34].
@@
Recently, co-simulation of SUMO and MetaDrive was used to train different RL algorithms for TSC at a single intersection using the raw pixel values of the image as observation input to the learning mechanism [9]. This incurs a notable performance cost of training on high dimensional data but allows setting up different visual conditions, thus reducing the simulation-reality gap. While MetaDrive can execute the simulation faster than real-time, the main drawback is that it lacks visual fidelity, so pixel values will be quite different in the real-world.
@@
Real-time video from 14 traffic cameras was used for vehicle detection and tracking using transfer learning, and the extracted information was incorporated into a simulation for TSC optimization [35]. The optimization was performed in simulation after analyzing previous 15 minutes of video and claims to reduce 14.2% of average vehicle delays. In another example of computer vision based signal timing control, vehicles were detected from video and represented in temporal clusters which were used to estimate road occupancy instead of counting [36]. Phase durations were determined from Gaussian regression, and while the inference process was fast enough for real-time execution, the framework was only valid for a single intersection. Similarly in [43], YOLOv4 was used to count vehicles and Greenshield’s model was used to adjust the actuation of the green time for a single intersection. The YOLO family of convolutional neural network based models are some of the most popular object detection models, designed to be fast and accurate with pretrained checkpoints available in various sizes [44].
@@
Reinforcement learning for TSC is generally performed using ground truth observations from the simulator itself, and the testing is also done in the same environment with varying traffic demands. A brief selection of papers involving co-simulation are compared in Table I with applications such as connected autonomous vehicles, traffic modeling, and traffic signal control. To the best of our knowledge, no simulation framework currently exists to support traffic signal control based on multi-agent reinforcement learning (MARL) using camera-based detection as input. This study addresses this gap by proposing and testing a visually realistic co-simulation environment that integrates video detection in CARLA with MARL signal control in SUMO.
III Methodology
III-A Overview

### v2

@@ Abstract
Traffic simulations are commonly used to optimize urban traffic flow, with reinforcement learning (RL) showing promising potential for automated traffic signal control, particularly in intelligent transportation systems involving connected automated vehicles. Multi-agent reinforcement learning (MARL) is particularly effective for learning control strategies for traffic lights in a network using iterative simulations. However, existing methods often assume perfect vehicle detection, which overlooks real-world limitations related to infrastructure availability and sensor reliability. This study proposes a co-simulation framework integrating CARLA and SUMO, which combines high-fidelity 3D modeling with large-scale traffic flow simulation. Cameras mounted on traffic light poles within the CARLA environment use a YOLO-based computer vision system to detect and count vehicles, providing real-time traffic data as input for adaptive signal control in SUMO. MARL agents trained with four different reward structures leverage this visual feedback to optimize signal timings and improve network-wide traffic flow. Experiments in a multi-intersection test-bed demonstrate the effectiveness of the proposed MARL approach in enhancing traffic conditions using real-time camera based detection. The framework also evaluates the robustness of MARL under faulty or sparse sensing and compares the performance of YOLOv5 and YOLOv8 for vehicle detection. Results show that while better accuracy improves performance, MARL agents can still achieve significant improvements with imperfect detection, demonstrating scalability and adaptability for real-world scenarios.
keywords:
@@
Computer Vision; Co-simulation Framework; Infrastructure Camera; Reinforcement Learning; Traffic Signal Control
articletype: Research Article
1 Introduction
@@
Traffic congestion is a significant issue in urban areas leading to increased travel times, fuel consumption, and environmental pollution, causing considerable economic impact (Goodwin, 2004; Luan et al., 2022; Qiu et al., 2020). Effective traffic signal control (TSC) can alleviate these problems by optimizing the flow of vehicles through intersections. Traditional TSC methods often rely on fixed-time schedules or simple rule-based systems, which may not adapt well to dynamic and complex traffic conditions (Mousavi et al., 2017). In the Capital Region of New York State, most locally-owned (non-DOT) signals are pre-timed, and all DOT signals are actuated, but none are coordinated for network-wide optimization except a small number of local signals. As a result, there is growing interest in applying advanced techniques, such as reinforcement learning (RL), to develop adaptive and efficient traffic signal control systems.
@@
Evaluating traffic control strategies in simulation before real-world deployment is essential for intelligent transportation systems (ITS) (Feng et al., 2023). Among the available simulation tools, Simulation of Urban MObility (SUMO) (Krajzewicz et al., 2002) stands out for its open-source availability, ease of use, and effectiveness (Li et al., 2024). While this is a sufficient and widely used environment for traffic signal control experimentation, the lack of realistic synthetic sensor data limits the applicability to real-world situations. Despite the widespread use of video cameras due to their low cost and ability to capture rich data (Ke, 2020), previous studies have yet to fully integrate them into SUMO-based traffic control systems. To bridge this gap, co-simulation frameworks that combine multiple simulators are gaining popularity as a means to create more realistic and comprehensive evaluation environments that include virtual cameras (He et al., 2024). However, further work is needed to transform these cameras into smart sensors that can interact with microsimulation control algorithms in a synchronized way. This integration would create a virtual test-bed with significant potential to simulate surveillance camera-based ITS control strategies.
@@
In this study, we propose and evaluate such a promising test-bed by exploring the use of multi-agent reinforcement learning (MARL) for traffic signal optimization and using camera-based detection for testing. In order to examine the effectiveness of the co-simulation framework, we develop and evaluate MARL agents in this realistic test-bed that integrates the CARLA simulator (Dosovitskiy et al., 2017) and the SUMO traffic simulator. CARLA is an open-source urban driving simulator that provides high-fidelity 3D environments for vehicle dynamics and sensor simulations, while SUMO is an open-source, highly portable, microscopic and continuous multi-modal traffic simulation package designed to handle large road networks (Ke, Weidner, et al., 2023).
@@
MARL involves training multiple RL agents, each controlling a traffic signal, to collectively improve overall travel time in the road network. This approach allows each traffic signal to adapt its timings based on real-time traffic conditions, leading to more efficient traffic management (X. Wang et al., 2020). There are numerous ways to train these agents in terms of architecture and reward functions. However, the challenges of real-world applicability stemming from faulty or sparse sensing, communication delays, and computational bottlenecks can significantly impact the effectiveness of any control system (Azfar et al., 2024), yet these issues are rarely addressed in simulation studies. This consideration is directly relevant to the application of connected automated vehicles (CAV), which depend on infrastructure-based sensing and communication. Demonstrating traffic control strategies that are resilient to noisy detection is therefore essential for supporting CAV deployment in practice.
@@
In our framework, CARLA is used to provide a realistic urban environment and detailed vehicle dynamics, while SUMO handles traffic flow and signal control. Cameras are deployed on traffic light poles within the CARLA environment to monitor traffic at intersections. In the Capital Region of New York State, most signals are on span wire, so cameras need to go on the poles. These cameras feed a convolutional neural network based computer vision system to detect and count vehicles, providing real-time data on traffic conditions which is used by the MARL agents to control the next phase of the traffic light. Figure 1 showcases the framework for the system. The agents have been trained using a Q-learning algorithm to optimize signal timings based on the observed traffic conditions. The goal is to minimize network-wide vehicle waiting times and improve overall traffic flow. This is the most common measure of effectiveness (MoE), while other MoEs like number of arrivals on red could also be used because it correlates with quality of traveler experience and vehicle emissions.
@@
This paper presents the methodology and results of our co-simulation study, demonstrating the potential of MARL for adaptive TSC using computer vision based detection that is often imperfect. We discuss the relevant literature followed by the setup of the CARLA-SUMO co-simulation environment, the implementation of the vision system for vehicle detection, and the training and evaluation of the MARL agents. Our results show that MARL-based TSC can significantly improve traffic flow compared to traditional fixed-time or actuated signal control methods while using inputs from infrastructure camera-based vehicle detection. We also discuss the errors related to detection and its effects on the system performance. In summary, the primary contributions of the paper are as follows:
@@
  • The framework programmatically accesses traffic light pole locations in the CARLA environment, spawns virtual cameras at these locations, and uses coordinate matching between CARLA and SUMO to identify the stopping line of the lane each light serves. It then automatically calculates the camera’s yaw angle to ensure it is properly aimed. These cameras use a YOLO-based computer vision system to detect and count vehicles, providing real-time data on traffic conditions in the virtual test-bed.
  • The study employs MARL with four different well known rewards to train multiple agents for adaptive traffic signal control, optimizing the timings based on feedback of real-time traffic conditions to improve network-wide traffic flow. Experiments on the co-simulation test-bed show that the proposed MARL significantly improves traffic signal control using camera-based detection.
  • The MARL agents are in scenarios with potentially faulty or sparse sensing, highlighting the importance of robustness in real-world applications. The paper examines performance differences between YOLOv5 and the newer YOLOv8, and how the detection errors affect the agents trained using different rewards, finding that while increased accuracy improves performance, even imperfect detection can yield positive results with the right MARL training.
@@ 2 Literature Review
There are numerous traffic simulation software available for different purposes including traffic flow modeling, simulating connected autonomous vehicles, vehicle dynamics and emissions modeling, and traffic signal control (Nguyen et al., 2021). Among the most widely used traffic simulators are AIMSUN (Casas et al., 2010), VISSIM (Fellendorf & Vortisch, 2010), and MATSim (Balmer et al., 2009), each offering unique features and capabilities to suit different research needs (Diallo et al., 2021). However, the focus of this study is on SUMO (Krajzewicz et al., 2002), a highly regarded open-source traffic simulator. The open-source nature of SUMO and the Traffic Control Interface (TraCI) enable finely customizable real-time programmable interaction with the simulation. It has a rich history of use in intelligent traffic signal control for example, Azevedo et al. (2016) created a toolchain to access and control SUMO traffic lights, enabling machine learning via external programming, and Guo et al. (2019) developed a Q-network for a single intersection in SUMO.
@@
While automated traffic signal control has been around for decades, most notably the Sydney coordinated adaptive traffic system (SCATS) (Sims & Dobinson, 1980) and split, cycle, and offset optimisation technique (SCOOT) (Hunt, 1981). These methods utilized real-time traffic flow data at each intersection to manage the timing of signal phases. The SCATS system uses vehicle counts at one intersection whereas SCOOT relies on upstream detectors for more detailed flow information. Both systems are capable of coordination with neighboring intersections by communicating their state information. However, the systems operate in a local context, their behavior is limited by the preprogrammed algorithm, and they are sensitive to the position and reliability of the vehicle sensors (Reza et al., 2023).
@@
Machine learning strategies, in particular reinforcement learning (RL) has emerged as an effective strategy for automated traffic control (M. Zhu et al., 2020). Coordinated deep Q-network RL was used with transfer planning for fast and scalable learning, but the training was found to occasionally lack in stability (Van der Pol & Oliehoek, 2016). Reinforced Signal Control (RESCO) (Ault and Sharon, 2021) provided a benchmark test-bed for RL based TSC inspired by real world cities and calibrated demand. Their extensive experimentation showed that decentralized control algorithms are more robust as compared to coordinated control methods, which are hampered by stricter sensing requirements and do not generalize well to all traffic conditions. A scalable multi-agent RL (MARL) algorithm for TSC was developed in Chu et al. (2019) which features independent advantage actor critic RL agents, each of which is kept updated about its neighbor’s policy. It was tested on a synthetic traffic grid and Monaco city traffic network, with the salient limitations being a lack of robustness to noisy or delayed sensor measurements, high memory consumption of the agent models, and communication latency between neighboring intersections. Xiao et al. (2022) introduced a staged dueling double deep Q-network integrated with model based control to avoid the cold start problem for RL. A strategy to simplify the optimization process in Xu et al. (2020) identifies the critical nodes using trajectory based CRRank analysis and only applies LSTM based RL for network-wide control. Asynchronous n-step Q-learning was used to train an adaptive TSC reducing delays compared to actuated control in Genders & Razavi (2019), but increased delays for left turns indicated the need to fine-tune the reward function.
@@
Regarding the interpretability of intelligent traffic control agents, it is demonstrated in Ault et al. (2020) that a regulatable control policy is one in which changes to the signal assignment can be understood as a direct consequence of changes in the state of the traffic. This is important for safety and regulatory considerations and is comparable to the actuated signal technology currently deployed based on government standards (Bonneson et al., 2011).
@@
A traffic simulator such as SUMO can perform its role effectively but is limited in its ability to provide synthetic sensor data beyond loop detectors. Additional simulation tools are required for problems that require real-world considerations such as demand prediction, advanced sensing, visual accuracy, vehicle mechanics, etc. These simulations must be synchronized for consistency and validity, and this process is known as co-simulation. MATSim was integrated with external simulators for fleet demand simulation in H. Yang et al. (2024) to showcase ride-pooling services with agent based simulation.
CARLA is an open source simulator for autonomous driving built atop the Unreal 4 game engine (Dosovitskiy et al., 2017). It features high quality 3D graphics, a flexible API, and integration with SUMO. CARLA-SUMO co-simulation is extensively used in many applications including the study of connected and autonomous vehicles (Stepanyants & Romanov, 2023), real-time sensor fusion (Cantas & Guvenc, 2021), computer vision based cooperative perception (Aoki et al., 2020), and transportation digital twins (Azfar et al., 2022). The co-simulation has also been used in conjunction with image processing and object detection to estimate traffic flow statistics such as volume and trajectories from real world video in order to generate simulation scenarios based on real-world conditions (Shi et al., 2022).
@@
Recently, co-simulation of SUMO and MetaDrive was used to train different RL algorithms for TSC at a single intersection using the raw pixel values of the image as observation input to the learning mechanism (He et al., 2024). This incurs a notable performance cost of training on high dimensional data but allows setting up different visual conditions, thus reducing the simulation-reality gap. Frameworks like MetaDrive+SUMO, while valuable, lack the visual fidelity required to test the robustness of vision algorithms against realistic environmental conditions. Our use of CARLA provides high-quality 3D graphics that allow for testing under varied weather and lighting, which is crucial for evaluating performance in real-world scenarios.
@@
Real-time video from 14 traffic cameras was used for vehicle detection and tracking using transfer learning, and the extracted information was incorporated into a simulation for TSC optimization (Zhou et al., 2021). The optimization was performed in simulation after analyzing previous 15 minutes of video and claims to reduce 14.2% of average vehicle delays. In another example of computer vision based signal timing control, vehicles were detected from video and represented in temporal clusters which were used to estimate road occupancy instead of counting (Kumaran et al., 2019). Phase durations were determined from Gaussian regression, and while the inference process was fast enough for real-time execution, the framework was only valid for a single intersection. Similarly in Gorodokin et al. (2021), YOLOv4 was used to count vehicles and Greenshield’s model was used to adjust the actuation of the green time for a single intersection. The YOLO family of convolutional neural network based models are some of the most popular object detection models, designed to be fast and accurate with pretrained checkpoints available in various sizes (Jocher, 2020). Aerial images from the VISSIM simulator were used to train RL-based signal control for a single intersection in Jeon et al. (2018), providing an inconsistent but significant reduction in delays.
Other approaches incorporate real-time information through probe vehicles (Lian et al., 2021) however, with limited penetration rates, their usefulness is largely restricted to congested conditions, while under normal flow they serve mainly as proxies for average speed. Multi-agent optimization based on dynamic programming and optimal control demonstrate superior performance but assume perfect communication and observations (Xu et al., 2019). The use of cellular data to estimate traffic flows is investigated in F. Yang et al. (2016) by including a realistic simulation of wireless signals in a traffic environment. However, in practice, traffic cameras are already widely deployed for monitoring and incident detection, and have been shown to be highly cost-effective, returning up to \$12 in benefits for every dollar spent (Fries et al., 2007).
@@
Reinforcement learning for TSC is generally performed using ground truth observations from the simulator itself, and the testing is also done in the same environment with varying traffic demands. A brief selection of papers involving co-simulation are compared in Table 1 with applications such as connected autonomous vehicles, traffic modeling, and traffic signal control.
Table 1: Existing literature related to traffic co-simulation
Paper Task Key Method Roadside Sensing   Scale Feedback Loop Visual Realism Osiński et al. (2020) RL for autonomous driving Image segmentation + CNN No Single vehicle Yes Yes Cantas & Guvenc (2021) Collaborative perception for CAVs LiDAR 3D detection + Sensor fusion No Multiple vehicles Yes Yes Aoki et al. (2020) Cooperative perception and V2X communication Object classification + Deep-Q RL No Multiple vehicles No Yes Azfar et al. (2023) Incorporating vehicle detection in Digital twin Edge-based vehicle detection Yes Single roundabout No Real video Shi et al. (2022) Data-driven test-bed for traffic scenes Vehicle detection and trajectory extraction Yes Highway segment or intersection No Real video Zhou et al. (2021) CV based traffic modeling and adaptive signal schemes Transfer learning for vehicle detection and tracking Yes Road network No No Kumaran et al. (2019) CV based traffic signal optimization Optical flow and clustering + Gaussian regression Yes Single intersection No Real video He et al. (2024) CV based isolated TSC End-to-end deep RL from images Yes Single intersection Yes No Zhao et al. (2022) Modeling and evaluating CAV and human behavior Mixed traffic modeling and game theory No Highway on-ramp Yes Yes Insuasty et al. (2019) Control strategies for urban traffic with co-simulation Particle swarm optimization + Replicator dynamics No Two intersections Yes No Y. Zhu et al. (2024) Mixed traffic stability analysis in co-simulation Mathematical modeling with V2V and car following No Multiple vehicles No No Jeon et al. (2018) Traffic signal control from VISSIM aerial images Deep CNN for Q function approximation No Single intersection No Real video Ours Test-bed for multi-agent traffic signal optimization using infrastructure camera input Vehicle detection + MARL Yes Multiple intersections Yes Yes
To the best of our knowledge, this is the first study to explicitly model and evaluate a network of infrastructure-mounted cameras for multi-agent reinforcement learning (MARL)-based traffic signal control. No existing simulation framework currently supports MARL-based signal control using realistically rendered, camera-based detection. To address this gap, we propose and demonstrate a visually realistic co-simulation environment that integrates video detection in CARLA with signal control in SUMO. The core of our evaluation is testing MARL robustness to imperfect sensing from realistic, ground-level cameras. CARLA runs on Unreal Engine and boasts realistic dynamic lighting, weather, and reflections. This cannot be replicated in low-fidelity simulators like MetaDrive, which trades off realistic visual rendering for performance. Other studies that utilize aerial imagery with VISSIM simplify the problem by adopting a bird’s-eye view, thereby bypassing the complexities associated with modeling ground-level camera infrastructure. Our framework is uniquely positioned to close the ‘simulation-reality gap’ for this specific problem.
3 Methodology
3.1 Overview